Consider the function 𝑦 = tan (π‘₯)


a) Show that the first two non-zero terms in the Maclaurin series of 𝑦 are "\ud835\udc65 + 1\/3 \ud835\udc65 ^3" …


b) Use the first two terms of the Maclaurin series of 𝑦 to estimate "tan ( 1\/3 )"


1
Expert's answer
2021-09-20T01:41:19-0400

"(i) Let \ f(x)=\tan x, then \ f(0)=0 \\
f^{\prime}(x)=\sec ^{2} x=1+\tan ^{2} x \\
 \Rightarrow f^{\prime}(0)=1 \\
f^{\prime \prime}(x)=2 \sec x\left(1+\tan ^{2} x\right)\\
 \Rightarrow f^{\prime \prime}(0)=2 \\

The \ Maclaurin \ series \ expansion \ is \\
 
\tan x=f(0)+x f^{\prime}(0)+\frac{x^{2}}{2 !} f^{\prime \prime}(x)+\frac{x^{3}}{3 !} f^{\prime \prime}(x)+\ldots \\
=x+\frac{x^{3}}{3}+o\left(x^{3}\right)"

"(ii) Put \ x=\frac{1}{3}\\
\therefore tan(\frac{1}{3})=\frac{1}{3}+\frac{1}{3}(\frac{1}{3})^3+...\\
=\frac{1}{3}+\frac{1}{81}\\
=\frac{28}{81}"